# Learning Hidden Physics and System Parameters with Deep Operator Networks

### v1

@@ Abstract
Big data is transforming scientific progress by enabling the discovery of novel models, enhancing existing frameworks, and facilitating precise uncertainty quantification, while advancements in scientific machine learning complement this by providing powerful tools to solve inverse problems to identify the complex systems where traditional methods falter due to sparse or noisy data. We introduce two innovative neural operator frameworks tailored for discovering hidden physics and identifying unknown system parameters from sparse measurements. The first framework integrates a popular neural operator, DeepONet, and a physics-informed neural network to capture the relationship between sparse data and the underlying physics, enabling the accurate discovery of a family of governing equations. The second framework focuses on system parameter identification, leveraging a DeepONet pre-trained on sparse sensor measurements to initialize a physics-constrained inverse model. Both frameworks excel in handling limited data and preserving physical consistency. Benchmarking on the Burgers’ equation and reaction-diffusion system demonstrates state-of-the-art performance, achieving average $L_{2}$ errors of $\mathcal{O}(10^{-2})$ for hidden physics discovery and absolute errors of $\mathcal{O}(10^{-3})$ for parameter identification. These results underscore the frameworks’ robustness, efficiency, and potential for solving complex scientific problems with minimal observational data.
@@
physics deficient equation, deep operator network, system identification and generalization, scientific machine learning, inverse problem.
@@ 1 Introduction
Recent advances in machine learning, coupled with innovative data recording and sensor technologies, are revolutionizing our understanding of complex systems across scientific and engineering domains. Traditional approaches often struggle to capture the intricate dynamics of systems in fields like neuroscience, epidemiology, and materials science, where direct measurement or explicit modeling proves challenging. The integration of statistical learning techniques with classical applied mathematics has emerged as a powerful approach to bridging this knowledge gap. Consider, for example, the challenge of understanding fluid dynamics [1] in aerospace engineering or predicting cardiovascular blood flow in medical research [2, 3] or predicting failure in materials [4]. These systems are characterized by nonlinear interactions and multiple interdependent variables that defy simple, first-principles modeling. Machine learning offers a transformative solution, enabling researchers to develop sophisticated mathematical models that directly emerge from empirical data. The landscape of data-driven dynamical systems discovery is rich and multifaceted, encompassing diverse methodological innovations. Researchers have developed approaches such as Equation-free modeling [5], deep learning techniques [6], non-linear regression [7], empirical dynamic modeling [8], automated dynamics inference [9], symbolic regression [10], and Koopman analysis [11] to address the fundamental challenge of modeling complex systems with incomplete knowledge and limited observational data.
@@
A critical frontier in this domain is discovering closed-form mathematical models of real-world systems described by partial differential equations (PDEs) and identifying crucial system parameters from sparse, spatiotemporally scattered data. Traditional approaches have struggled to bridge the gap between theoretical understanding and practical measurement. Several prominent approaches have emerged to address these challenges. Existing methods like Sparse Identification of Non-linear Dynamics [12] (SINDy) and Deep Hidden Physics Models [6] (DHPM) have made significant strides, but they face inherent limitations. SINDy approximates system behavior using sparse linear combinations of predefined candidate terms, but the framework requires prior knowledge of system non-linearities and sensitivity to measurement noise. Deep Hidden Physics Models (DHPM) offers a promising alternative by utilizing two deep neural networks. One network captures the system state, while another approximates the unknown physics of the model. The DHPM framework demonstrates remarkable stability, leveraging automatic differentiation techniques to compute gradient terms, thus effectively bypassing the numerical differentiation challenges inherent in traditional approaches. Deep Hidden Physics Models (DHPM) offer improved stability through dual neural networks - one capturing system state, another approximating unknown physics - but cannot identify unknown system parameters from labeled datasets.
Physics-Informed Neural Networks (PINNs) [13, 14, 15] marked a breakthrough in parameter estimation by introducing innovative functional approximation techniques. These methods simultaneously minimize both the residual loss from governing PDEs and the data loss from sparse sensor recordings, demonstrating success across diverse applications - from estimating hemodynamic parameters in cardiovascular systems [16] to reconstructing elasticity fields in heterogeneous materials [4], as well as inferring flow parameters in fluid mechanics [17, 18]. However, a fundamental limitation persists: existing methods typically require re-training and re-evaluation for each system parameter variation, limiting their generalizability and computational efficiency.
This persistent challenge motivates the need for more robust and adaptable modeling techniques. Motivated by these challenges, this study introduces two complementary frameworks:
@@
    The Deep Hidden Physics Operator (DHPO) network - a generalized operator identified from minimal labeled datasets spanning various system conditions. This framework uniquely integrates DHPM [6] with neural operators [19], specifically deep operator networks, to model system dynamics with partially known physics.
@@
    A physics-informed operator-learning framework that identifies system parameters for sparse sensor recordings across PDE families, requiring no additional labeled dataset and operating in a semi-supervised approach.
Both these frameworks build upon recent advances in neural operators [20] designed to learn the mappings between infinite dimensional functional spaces, thus facilitating generalization abilities. There are numerous neural operators, which can be categorized into meta-architectures, such as deep operator networks (DeepONet) [19], and operators based on integral transforms, including the Fourier neural operator (FNO) [21], wavelet neural operator (WNO) [22], graph kernel network (GKN) [23], convolutional neural operator [24], and Laplace neural operator (LNO) [25], basis-to-basis framework [26], and the resolution-independent neural operator [27] among others. Among the various neural operators developed, in this work, we consider the DeepONet for its architectural flexibility. Its architecture, inspired by the universal approximation theorem for operators [28], consists of a branch network that encodes varying input functions and a trunk network that encodes spatio-temporal coordinates. Since its first appearance, vanilla DeepONet architecture has been employed to tackle challenging problems involving complex high-dimensional dynamical systems [29, 30, 31, 32]. In addition, extensions of DeepONet have been recently proposed in the context of integration of multiple-input continuous operators [33, 34], hybrid transferable numerical solvers [35, 36], transfer learning [37], physics-informed (PI) learning to satisfy the underlying PDE [38, 20, 39], and multi-task learning framework [40]. Recent advancements have significantly improved its capabilities, including efficient training strategies [41], Seperable PI-DeepONet’s efficient PDE solving [39], RI-DeepONet for resolution-independent training [27], Geom-DeepONet’s 3D geometry prediction [42], and L-DeepONet’s for learning operators in latent space mapping [43]. These developments have enhanced computational efficiency and predictive accuracy across diverse scientific domains. In this study, we employ the vanilla DeepONet as our foundational framework.
These advancements represent a significant step toward robust and efficient modeling of complex dynamical systems with limited observational data. By pushing the boundaries of data-driven scientific modeling, the proposed frameworks offer new opportunities for understanding and predicting intricate system behaviors across various scientific and engineering domains. The remainder of the paper is organized as follows. Section 2 outlines the framework for building surrogates for hidden physics discovery and parameter identification. Section 3 demonstrates the applicability of the proposed framework using benchmark examples. Section 4 summarizes our study and highlights potential future research directions.
@@ 2 Physics Discovery and Parameter Identification using Neural Operators
The successful application of deep learning in physical sciences hinges on two critical capabilities: discovering underlying physics from data and identifying system parameters that govern physical processes. Traditional approaches often struggle with generalization across different physical scenarios and require extensive data collection. In this section, we present two frameworks that address these challenges by leveraging DeepONet: first, an innovative approach for discovering physical laws that can generalize across multiple PDE systems, and second, an operator framework for system parameter identification that efficiently handles inverse problems in physical systems.
@@ 2.1 Deep Hidden Physics Operator (DHPO) - Discovering physics using DeepONet
Building upon the concept of deep hidden physics models (DHPM) introduced by Raissi et al. [6], we introduce deep hidden physics operator (DHPO) - a framework that leverages the power of operator learning to discover underlying physical laws. While DHPM has proven effective in learning physics from data corresponding to a single PDE system under varying boundaries or initial conditions, our approach extends this capability to learn from multiple PDE systems simultaneously.
@@
for $x\in\Omega,t\in[0,T],u:\Omega\times t\rightarrow\mathbb{R}$. Here, $\mathcal{N}$ represents the unknown physics operator we aim to discover, and $f(x)$ is a variable source term. Our framework, DHPO, employs a DeepONet architecture consisting of two primary components: a branch network that processes the input function $f(x)$ at fixed sensor locations, and a trunk network that handles the spatio-temporal coordinates $(x,t)$. The solution field $u(x)$ is approximated through the interaction of these networks via a dot product operation. The gradient, $u_{t},u_{x},u_{xx}$, is computed using the automatic differentiation approach. To discover the underlying physics, we introduce a hidden physics neural network that approximates $\mathcal{N}$ by taking the tuple $[u,u_{x},u_{xx}]$ as inputs. This network learns to represent the unknown terms in the PDE while maintaining the interpretability of the discovered physics. The training of this framework is governed by a composite loss function that encompasses:
@@
2.2 System Parameter Identification using DeepONet
The framework consists of two key components. First, a DeepONet is pre-trained to learn the forward mapping from sparse input data to the complete solution field. Second, we introduce an additional MLP designed to learn the inverse mapping from the solution space to the parameter space. These two networks – the DeepONet and MLP – are then trained through continual learning, using the PDE residual as the primary loss term. To ensure solution accuracy and physical consistency, we incorporate additional loss terms for known data.
1 Inputs: sensor measurements $f(x^{s})$ and $u(x^{s},t^{s})$ at $n$ fixed locations, collocation points $x,t$ , partially known physics $\mathcal{M}$, initial conditions, boundary conditions
@@ 11             $u_{\text{ic}}$ = $\mathcal{G}_{\bm{\theta}}(f(x^{s})(x,0))$
12             $u_{\text{bc}}$ = $\mathcal{G}_{\bm{\theta}}(f(x^{s})(x_{\text{bc}},t))$
@@ 17             Forward pass through MLP
18             $\mathcal{N}=\text{MLP}(u_{\text{pred}},u_{x_{\text{pred}}},u_{xx_{\text{pred}}},...)$
@@ 19            Compute loss
20             $\mathcal{L}^{\text{pde}}=\frac{1}{N}\sum_{i=1}^{N}||\mathcal{N}_{i}+\mathcal{M}_{i}-f(x^{s})_{i}||^{2}$   physics loss
21             $\mathcal{L}^{\text{ic}}=\frac{1}{N}\sum_{i=1}^{N}||u_{\text{ic}_{i}}-u(x,0)||^{2}$   initial condition loss
22             $\mathcal{L}^{\text{bc}}=\frac{1}{N}\sum_{i=1}^{N}||u_{\text{bc}_{i}}-u(x_{\text{bc}},t)||^{2}$   boundary condition loss
23             $\mathcal{L}^{\text{data}}=\frac{1}{N}\sum_{i=1}^{N}||u_{\text{pred}_{i}}-u(x^{s},t^{s})||^{2}$   data loss
24             $\mathcal{L}(\bm{\theta},\phi)=\mathcal{L}^{\text{pde}}(\bm{\theta},\phi)(x,t)+\mathcal{L}^{\text{ic}}(\bm{\theta})(x,0)+\mathcal{L}^{\text{bc}}(\bm{\theta})(x_{\text{bc}},t)+\mathcal{L}^{\text{data}}(\bm{\theta})(x^{s},t^{s})$
@@ 25             Backpropagate and update $\bm{\theta}$ and $\phi$ using Adam
1 Inputs: sensor measurements $u(x^{s},t^{s})$ at $n$ fixed locations, collocation points $x,t$ , PDE $\mathcal{N}(\mathcal{X},u(x,t),\xi)$

### v2

@@ Abstract
Discovering hidden physical laws and identifying governing system parameters from sparse observations are central challenges in computational science and engineering. Existing data-driven methods, such as physics-informed neural networks (PINNs) and sparse regression, are limited by their need for extensive retraining, sensitivity to noise, or inability to generalize across families of partial differential equations (PDEs). In this work, we introduce two complementary frameworks based on deep operator networks (DeepONet) to address these limitations. The first, termed the Deep Hidden Physics Operator (DHPO), extends hidden-physics modeling into the operator-learning paradigm, enabling the discovery of unknown PDE terms across diverse equation families by identifying the mapping of unknown physical operators. The second is a parameter identification framework that combines pretrained DeepONet with physics-informed inverse modeling to infer system parameters directly from sparse sensor data. We demonstrate the effectiveness of these approaches on benchmark problems, including the Reaction-Diffusion system, Burgers’ equation, the 2D Heat equation, and 2D Helmholtz equation. Across all cases, the proposed methods achieve high accuracy, with relative solution errors on the order of $\mathcal{O}(10^{-2})$ and parameter estimation errors on the order of $\mathcal{O}(10^{-3})$, even under limited and noisy observations. By uniting operator learning with physics-informed modeling, this work offers a unified and data-efficient framework for physics discovery and parameter identification, paving the way for robust inverse modeling in complex dynamical systems.
@@
physics-deficient equation, deep operator network, system identification and generalization, scientific machine learning, inverse problem.
@@ 1 Introduction
Recent advances in machine learning, coupled with increasingly sophisticated data acquisition technologies, are transforming how we model and understand complex physical systems. In domains such as fracture analysis in electromechanical systems Tarafder et al. [2023], fluid dynamics Perdikaris and Karniadakis [2016] in aerospace engineering, or predicting cardiovascular blood flow in medical research Kennedy et al. [2017], Gennisson et al. [2013], and failure in materials Chen and Gu [2021], classical first-principles models often struggle to capture nonlinear dynamics or account for incomplete knowledge of governing processes. Sparse or noisy observations further challenge conventional numerical and statistical methods. As a result, there has been a growing interest in integrating machine learning with applied mathematics to bridge these gaps and accelerate scientific discovery. The landscape of data-driven dynamical systems discovery is rich and multifaceted, encompassing diverse methodological innovations. Researchers have developed approaches such as Equation-free modeling Kevrekidis et al. [2004], deep learning techniques Raissi [2018], non-linear regression Motulsky and Ransnas [1987], empirical dynamic modeling Chang et al. [2017], automated dynamics inference Daniels and Nemenman [2015], symbolic regression Koza [1992], and Koopman analysis Brunton et al. [2021] to address the fundamental challenge of modeling complex systems with incomplete knowledge and limited observational data.
@@
A critical frontier in this domain is discovering closed-form mathematical models of real-world systems described by partial differential equations (PDEs) and identifying crucial system parameters from sparse, spatio-temporally scattered data. Traditional approaches have struggled to bridge the gap between theoretical understanding and practical measurement. Several prominent approaches have emerged to address these challenges. Existing methods like Sparse Identification of Non-linear Dynamics Brunton et al. [2016] (SINDy) and Deep Hidden Physics Models Raissi [2018] (DHPM) have made significant strides, but they face inherent limitations. SINDy approximates system behavior using sparse linear combinations of predefined candidate terms, but the framework requires prior knowledge of system non-linearities and is sensitive to measurement noise as well as is typically designed for recovering explicit governing equations for a single dynamical system. Deep Hidden Physics Models (DHPM) offers a promising alternative by utilizing two deep neural networks. One network captures the system state, while another approximates the unknown physics of the model. The DHPM framework demonstrates remarkable stability, leveraging automatic differentiation techniques to compute gradient terms, thus effectively bypassing the numerical differentiation challenges inherent in traditional approaches. DHPM offers improved stability through dual neural networks - one capturing system state, another approximating unknown physics - but cannot identify unknown system parameters from labeled datasets. Physics-Informed Neural Networks (PINNs) Raissi et al. [2019], Cai et al. [2021], Wang et al. [2024] marked a breakthrough in parameter estimation by introducing innovative functional approximation techniques. These methods simultaneously minimize both the residual loss from governing PDEs and the data loss from sparse sensor recordings, demonstrating success across diverse applications - from estimating hemodynamic parameters in cardiovascular systems Kissas et al. [2020] to reconstructing elasticity fields in heterogeneous materials Chen and Gu [2021], as well as inferring flow parameters in fluid mechanics Raissi et al. [2020], Zhang et al. [2023]. However, a fundamental limitation persists: existing methods typically require re-training and re-evaluation for each system parameter variation, limiting their generalizability and computational efficiency.
These challenges motivate the development of more flexible frameworks that (i) generalize across families of PDEs, (ii) infer hidden physics and system parameters simultaneously, and (iii) remain robust under sparse and noisy observations. Neural operators Lu et al. [2021], Goswami et al. [2022], Li et al. [2021], Tripura and Chakraborty [2023], Raonic et al. [2023], Cao et al. [2024] offer a compelling solution by learning mappings between infinite-dimensional functional spaces, thereby enabling operator-level generalization across different initial conditions, boundary conditions, or source terms. Among these, deep operator networks (DeepONet) Lu et al. [2021] are particularly appealing due to their architectural flexibility and proven success in modeling high-dimensional dynamical systems Kontolati et al. [2024], Agarwal et al. [2025], Wang et al. [2025], Sarkar et al. [2025b]. Recently, Molinaro et al. Molinaro et al. [2023] proposed a Neural Inverse Operator (NIO) methodology that learns operator-to-function mappings for canonical inverse problems, including Electrical Impedance Tomography, wave scattering, and seismic imaging. Their approach leverages multiple boundary measurements to characterize the observation operator and requires paired datasets of measurements and ground truth parameters generated from numerical simulations. Similarly, Long et al. Long et al. [2025] introduced an invertible Fourier Neural Operator (iFNO) that jointly learns forward and inverse mappings through a reversible architecture combined with variational autoencoders, enabling bidirectional predictions given complete paired training data. While these methods demonstrate the potential of neural operators for inverse problems, they fundamentally rely on the availability of labeled parameter datasets or complete field observations for training. In contrast, our work addresses scenarios where ground truth parameters are unavailable and only sparse sensor measurements are accessible, leveraging physics-informed constraints to enable parameter discovery directly from governing equations rather than paired supervision. In this study, we introduce two complementary physics-informed operator learning frameworks, built upon recent advances in neural operators:
@@
    The Deep Hidden Physics Operator (DHPO) network: The framework integrates DHPM with DeepONet to discover unknown governing physics directly from sparse data. Unlike DHPM, DHPO generalizes across multiple PDE families and system conditions, while maintaining interpretability by approximating the hidden terms of the governing equations.
@@
    Physics-Informed Parameter Identification: A framework for inferring system parameters (e.g., viscosity, diffusivity, thermal conductivity, source function) using sparse sensor recordings. Here, a pretrained DeepONet reconstructs the solution field from limited data, which then serves as input to a physics-constrained inverse model for parameter estimation. Unlike data-driven inverse operator approaches that require labeled parameter datasets, our framework addresses the more challenging scenario where ground truth parameters are unavailable during training. We leverage the known physics of the governing equations as constraints, enabling parameter identification purely from sparse sensor observations and PDE residuals without requiring supervised parameter labels.
By uniting operator learning with physics-informed modeling, this work contributes a step toward generalizable, data-efficient solvers for inverse problems in scientific computing. The proposed frameworks provide new opportunities for tackling hidden-physics discovery and parameter inference across diverse applications in engineering and the physical sciences. Both frameworks achieve high accuracy with limited observations, delivering relative $L_{2}$ errors on the order of $10^{-2}$ for hidden-physics discovery and absolute errors on the order of $10^{-3}$ for parameter identification. We demonstrate their performance on benchmark problems including the Burgers’ equation, a Reaction–Diffusion system, the 2D Heat equation, and the 2D Helmholtz equation and showing that our methods remain accurate and robust even under out-of-distribution scenarios. The remainder of the paper is organized as follows. Section 2 outlines the framework for building surrogates for hidden physics discovery and parameter identification. Section 3 demonstrates the applicability of the proposed framework using benchmark examples. Section 4 summarizes our study and highlights potential future research directions.
@@ 2 Physics Discovery and Parameter Identification using Neural Operators
The application of deep learning to the physical sciences hinges on two essential capabilities: uncovering governing equations directly from data when the underlying physics is only partially known, and estimating system parameters that control the dynamics of complex processes. Conventional data-driven approaches often fall short in these tasks, as they require large amounts of training data and generalize poorly across varying boundary conditions, source terms, or material properties. To address these limitations, we propose two complementary operator-learning frameworks based on DeepONet. The first, the Deep Hidden Physics Operator (DHPO), extends hidden-physics modeling into the operator-learning paradigm, enabling the discovery of unknown PDE terms across multiple equation families. The second is a parameter identification framework that leverages pretrained DeepONet to reconstruct solution fields from sparse sensor measurements and infer system parameters in a physics-consistent manner. Together, these approaches provide a unified, data-efficient strategy for physics discovery and parameter estimation in complex dynamical systems.
@@ 2.1 Deep Hidden Physics Operator (DHPO) - Discovering physics using DeepONet
Building upon the concept of deep hidden physics models (DHPM) introduced by Raissi et al. Raissi [2018], we introduce deep hidden physics operator (DHPO) - a framework that leverages the power of operator learning to discover underlying physical laws by identifying the mapping of unknown physical operators. While DHPM has proven effective in learning physics from data corresponding to a single PDE system under varying boundaries or initial conditions, our approach extends this capability to learn from multiple PDE systems simultaneously.
@@
for $x\in\Omega,t\in[0,T],u:\Omega\times t\rightarrow\mathbb{R}$. Here, $\mathcal{N}$ represents the unknown physics operator we aim to discover, and $f(x)$ is a variable source term. Our framework, DHPO, employs a DeepONet architecture consisting of two primary components: a branch network that processes the input function $f(x)$ at fixed sensor locations, and a trunk network that handles the spatio-temporal coordinates $(x,t)$. The solution field $u(x)$ is approximated through the interaction of these networks via a dot product operation. The gradient, $u_{t},u_{x},u_{xx}$, is computed using the automatic differentiation approach. To discover the underlying physics, we introduce a hidden physics neural network that approximates $\mathcal{N}$ by taking the tuple $u,u_{x},u_{xx}]$ as inputs. This network learns to represent the unknown terms in the PDE while maintaining the interpretability of the discovered physics. It is important to note that interpretability here does not refer to symbolic equation recovery, but to mechanistic interpretability: identifying which physical processes are active, their relative contributions, and their functional dependence on solution gradients. The training of this framework is governed by a composite loss function that encompasses:
@@
1Inputs: sensor measurements $f(x^{s})$ and $u(x^{s^{\prime}},t^{s^{\prime}})$ at $n$ fixed locations, collocation points $x,t$ , partially known physics $\mathcal{M}$, initial conditions, boundary conditions. $f(x)$ can be IC/BC/source or forcing $g(x)$.
@@ 11    $u_{\text{ic}}$ = $\mathcal{G}_{\bm{\theta}}(f(x^{s})(x,0))$
12    $u_{\text{bc}}$ = $\mathcal{G}_{\bm{\theta}}(f(x^{s})(x,t))$
@@ 17    Forward pass through MLP
[…]
@@ 19    Compute loss
[…]
24    $\mathcal{L}(\bm{\theta},\phi)=\mathcal{L}^{\text{pde}}(\bm{\theta},\phi)(x,t)+\mathcal{L}^{\text{ic}}(\bm{\theta})(x,0)+\mathcal{L}^{\text{bc}}(\bm{\theta})(x,t)+\mathcal{L}^{\text{data}}(\bm{\theta})(x^{s^{\prime}},t^{s^{\prime}})$
@@ 25      Backpropagate and update $\bm{\theta}$ and $\phi$ using Adam
We have considered the loss tolerance $\epsilon$ of $\mathcal{O}(10^{-3})$. To measure the deviation of the predicted solution from reference we use metric relative $L_{2}$ error and mean absolute error.
$\text{Relative}\ L_{2}\ \text{Error}=\frac{\sqrt{\sum_{i=1}^{N}|u_{i}-u_{i}^{*}|^{2}}}{\sqrt{\sum_{i=1}^{N}|u_{i}^{*}|^{2}}},$ (7)
$\text{Mean Absolute Error}=\frac{1}{N_{s}}{\sum_{i=1}^{N_{s}}E_{i};\text{where}\ E_{i}=\sum_{k=1}^{N_{i}}\frac{|u_{i}^{k}-u_{i}^{*k}|}{N_{i}}},$ (8)
where $u^{*},\;u$ are the reference and the predicted solution respectively. $N_{i}$ is number of data points for $i^{th}$ sample, $N_{s}$ is the total number of samples. We also define, the error distribution by $\mu\pm\sigma$, where $\mu=mean\{E_{i}\},\sigma=std.dev\{E_{i}\}$ are mean and standard deviation of errors over the samples. Unless otherwise noted, the mean test error refers to the mean absolute error (MAE) on the test set.
2.2 System Parameter Identification using DeepONet
2.2.1 Deterministic Parameter Estimation
The framework consists of two key components. First, a DeepONet is pre-trained to learn the forward mapping from sparse input data to the complete solution field. Second, we introduce an additional MLP designed to learn the inverse mapping from the solution space to the parameter space. These two networks - the DeepONet and MLP - are then trained through continual learning, using the PDE residual as the primary loss term. To ensure solution accuracy and physical consistency, we incorporate additional loss terms for known data.
2.2.2 Probabilistic Parameter Estimation
The deterministic framework provides a single point estimate of the unknown parameter $\xi$. However, inverse problems often suffer from non-uniqueness where multiple parameter values can produce similar observations given sparse and noisy sensor data. To capture this inherent ambiguity, we extend our framework to estimate the full probability distribution of parameters. The MLP now outputs both the mean $\bm{\mu}_{\xi}$ and standard deviation $\bm{\sigma}_{\xi}$ of a Gaussian distribution: $\xi\sim\mathcal{N}(\bm{\mu}_{\xi},\text{diag}(\bm{\sigma}_{\xi}^{2}))$, where $\bm{\mu}_{\xi},\bm{\sigma}_{\xi}]=\mathcal{I}_{\phi}(\mathbf{u}^{s})$. During training, we employ the reparameterization trick ($\xi=\bm{\mu}_{\xi}+\bm{\sigma}_{\xi}\odot\bm{\epsilon}$ with $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$) to enable backpropagation through the stochastic sampling process. The total loss function includes an additional KL divergence regularization term: $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{pde}}(\bm{\theta},\bm{\phi})+\lambda_{\text{data}}\mathcal{L}_{\text{data}}(\bm{\theta})+\lambda_{\text{KL}}D_{\text{KL}}\left(\mathcal{N}(\bm{\mu}_{\xi},\text{diag}(\bm{\sigma}_{\xi}^{2}))\,\|\,p_{\text{prior}}(\xi)\right)$, where the KL term constrains the learned distribution to remain within physically meaningful bounds based on a prior $p_{\text{prior}}(\xi)$ informed by domain knowledge.
The probabilistic framework offers distinct advantages, it naturally quantifies epistemic uncertainty (narrow distributions indicate high confidence, broad distributions signal ambiguity), explicitly captures parameter non-uniqueness rather than forcing potentially misleading point estimates, accommodates measurement noise through learned variance that increases with data corruption, and enables risk-aware decision making through access to the full parameter distribution for uncertainty propagation. The training procedure follows Algorithm 2 with minor modifications: the MLP output dimension is doubled to predict both $\bm{\mu}_{\xi}$ and $\bm{\sigma}_{\xi}$, parameter sampling via reparameterization is performed before physics loss evaluation, and the KL divergence term is added to the total loss. In Section 3, we demonstrate both deterministic and probabilistic parameter identification, showing that the probabilistic framework provides calibrated uncertainty estimates that accurately reflect parameter ambiguity in sparse observation scenarios.
1Inputs and Training Data: sensor measurements $u(x^{s},t^{s})$ at $n$ fixed locations, collocation points $(x,t)$, PDE $\mathcal{N}(\mathcal{X},u(x,t),\xi)$, and training pairs $(u_{i}(x^{s},t^{s}),\xi_{i})_{i=1}^{N_{\text{train}}}$ consisting of $N_{\text{train}}$ system realizations with corresponding true parameters.